OpenStudy (anonymous):

simplify 2√216 what the heck do i do?? :P

OpenStudy (anonymous):

Well, you simplify it.

OpenStudy (anonymous):

See what perfect square goes into 216. 216 is not a perfect square, so work down. As it happens, 216 divides into 36, which is 6^2, a perfect square. So, you would get \[2\sqrt{36 *6}\] Then square root the 36 (that's what it has to be a perfect square!) and bring it to the front: \[2*6\sqrt{6}\] \[12\sqrt{6}\]

OpenStudy (anonymous):

thanks! i understnad!;)